• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jpchen

    Log entry time 13:25:50 on November 9,2009

    Entry number 298583

    This entry is a followup to: 298508

    keyword=Target: temperature flucation, cryo flow fluctuation, hp heater.

    Target temperature fluctated quite a bit during the owl shift.
    I talkd to Kai and monitored for some time. It appeared to be due to the cryo flow fluctuation and frequent beam trips. Target has been responding ok. Both PID and beam compensation are working. The PID parameters have been used for long time and should be reasonable. We could try to play with these parameters to try optimize more, but probably won;t help much. The main thing is the cryo fluctuation. Due to large Delta_T with 4K cooling, these fluctuation has a large effect on the temperature and heater power. Even with these relatively large temperature changes, seems that the local density fluctuation does not increase significantly (from lumi data early day shift). If this is indeed the case, then we can live with the relative large temperature fluctuations.
    I changed alarm point for temperature from 22.5 to 23K. Will continue monitor and make adjustment as needed.